OpenStudy (anonymous):

solve by completing the square: 3x^2-9x-12=0

OpenStudy (anonymous):

3x^2-9x-12=0 or3(x^2-3x-4)=0 x^2-3x-4=0 x^2-2(3/2)x +(3/2)^2 -(3/2)^2 -4=0 (x-3/2)^2 -(9/4+4)=0 (x-3/2)^2 -(5/2)^2=0 (x-3/2-5/2)(x-3/2+5/2)=0 (x-4)(x+1)=0 or (x-4)=0 or (x+1)=0 hence x=4,-1

OpenStudy (anonymous):

thanks!!

OpenStudy (anonymous):

welcome dear